OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

simply rearrange for y = ax + b M = 2x + 3y 2x + 3y = M 3y = -2x + M y = -2/3x + M/3

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem.